OpenStudy (anonymous):

REWARD FOR HELP!!! A standard number cube is tossed. Find the probability. P(1 or odd)

OpenStudy (experimentx):

3/6 = 1/2

myininaya (myininaya):

\[P(1 or odd)=P(1)+P(odd)-P(1 and odd)=\frac{1}{6}+(P(1)+P(3)+P(5))-\frac{1}{6}=\frac{3}{6}\]